OpenStudy (anonymous):

the sum of three consecutive even integers is 348.find the sum of three

OpenStudy (anonymous):

x = 1st integer x + 2 = 2nd integer x + 4 = 3rd integer x + x + 2 + x + 4 = 348 now combine like terms and solve for x

OpenStudy (anonymous):

concentrate on the left side..what is x + x + x ?

OpenStudy (anonymous):

do you know ?

OpenStudy (anonymous):

hello ??